OpenStudy (anonymous):

Find the number of solutions using the quadratic formula; b^2 - 4ac. 5x^2 - 4x = -6

OpenStudy (anonymous):

First, you need to add 6 to both sides. 5x^2 - 4x +6 = 0

OpenStudy (anonymous):

In this equation a=5, b= -4, and c=6. Do you understand where I got those numbers?

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

but where do you add the 6 on both sides?

OpenStudy (anonymous):

5x^2 - 4x = -6 __________+6 __+6 5x^2 - 4x +6 = 0 In order to use the quadratic formula, one side of the equation has to be zero, so we add 6 to both sides of the equation to "move" it to the left side.

OpenStudy (anonymous):

Now that you have a, b, and c, you can plug them into the equation. If number you get is positive, there are 2 solutions, if it's negative, there aren't any real number solutions (I don't think you're working with imaginary numbers right now), and if the answer you get is 0, then there is one solution.

OpenStudy (anonymous):

okay thank you! :)

OpenStudy (precal):

b^2 - 4ac is know as the discriminant